Second Chance Welding rebuilds spot welders for manufacturers. The following budgeted cost data for 2017 is available for Second Chance. Time Charges Material Loading Charges Technicians' wages and benefits $237,000 - Parts manager's salary and benefits - $39,100 Office employee's salary and benefits 71,100 9,169 Other overhead 39,500 19,300 Total budgeted costs $347,600 $67,569 The company desires a $34.00 profit margin per hour of labor and a 21.00% profit margin on parts. It has budgeted for 7,900 hours of repair time in the coming year, and estimates that the total invoice cost of parts and materials in 2017 will be $386,000. Compute the rate charged per hour of labor. (Round answer to 2 decimal places, e.g. 10.50.) Labor rate $ per hour Compute the material loading percentage. (Round answer to 3 decimal places, e.g. 10.501.) Material loading percentage % Pace Corporation has requested an estimate to rebuild its spot welder. Second Chance estimates that it would require 41 hours of labor and $3,000 of parts. Compute the total estimated bill. (Round answer to 2 decimal places, e.g. 10.50.) Total estimated bill $
Answer #1
1
Budgeted cost per labor hour 44.00 =347600/7900
Add: Profit margin 34.00
Labor rate 78.00 per hour
2
Material loading charges 17.505% =67569/386000
Add: Profit on parts 21.00%
Material loading percentage 38.505%
3
Labor charges 3198.00 =41*78
Materials parts 3000.00
Material loading charges 1155.15 =3000*38.505%
Total estimated bill 7353.15